At what pressure will a sample of He (g) have a 

density of 0.112 g/L at 27°C?


1
Expert's answer
2021-02-03T07:24:45-0500

PM = "\rho" RT

M (molar mass of helium) = 4

R = 0.0821

T = 273+27 = 300K

"\rho" = 0.112g/L

P×4 = 0.112×0.0821×300

P = 0.68964 atm